ROTATIONAL TEMPERATURE OF METHANE IN SUPERSONIC MOLECULAR BEAM

\underline{JIAN TANG}, KEIKO SETOYAMA, TSUYOSHI HIRAO, AND NAOMASA UESUGI, Department of Chemistry, Faculty of Science, Okayama University, Tsushima-Naka 3-1-1, Okayama, Japan.

Rotational temperature of methane in supersonic molecular beam was claimed to be about 10 ∼ 40 K in various conditions. It was found that distribution of the three nuclear spin modifications do not change during the supersonic cooling. A strong deviation from thermal equillibrium was also observed. We re-investigated the rotational temperature of methane by the rovibrational transitions of the ν_{3} vibrational band. An infrared spectrometer with cw-OPO laser was used for the observation. The rotational temperature of CH_{4} derived from transitions up to J = 2 was 10 K, consistent with the rotational temperature derived for the OCS molecule. Much higher rotational temperature (∼ 70 K) with large diversity was obtained for higher J transitions of CH_{4}, which could be attributed to the incomplete relaxation of rotational-level distribution of light molecule as CH_{4}, a similar situation as the vibrational-level distribution in supersonic molecular beam.